OpenStudy (anonymous):

solve the system of equations using substitution 3y=-1/2x+2 y=-x+9

OpenStudy (xguardians):

Step One: Isolate a Variable of your choice in either of the equations Step Two: Plug in that variable. Step 3: Solve for the variable. 3(-x+9) = -1/2x + 2 -3x + 27 = -1/2 + 2 -2.5x = 25 x = -10 Step 3: Input the variable back into either of the equations. y = -(-10) + 9 y = 10 + 9 y = 19

OpenStudy (anonymous):

I dont have any choices as those answers

OpenStudy (anonymous):

@xGuardians

OpenStudy (xguardians):

Tell me your choices.

OpenStudy (anonymous):

they are (3,6) (-1,8) (20,-4) (10,-1)

OpenStudy (xguardians):

Bah messed up on my arithmetic haha, one second.

OpenStudy (anonymous):

no worries

OpenStudy (xguardians):

Step One: Isolate a Variable of your choice in either of the equations Step Two: Plug in that variable. Step 3: Solve for the variable. 3(-x+9) = -1/2x + 2 -3x + 27 = -1/2 + 2 -2.5x = -25 x = 10 Step 3: Input the variable back into either of the equations. y = -(10) + 9 y = -10 + 9 y = 1 (10,-1) is the answer, this is why you should always double check :P

OpenStudy (anonymous):

Thank you!